OpenStudy (anonymous):

check work: part a.) the area of a square is (16x^2-8x+1) square units. determine the length of each side of the square by factoring the areas expressions completely. show all work part b.) the area of a rectangle is (81x^2-4y^2)square units. determine the dimensions of the rectangle by factoring the area of the expression. show work c.) the volume of a rectangular box is (x^3-9x^2-4x+36) cubic units. determine the dimensions of the rectangular box by factoring the volume expression completely. show all work

OpenStudy (anonymous):

a. 16x^ 2 −4x−4x+1=4x(4x−1)−1(4x−1)=(4x−1)(4x−1)=(4x−1) 2 Length of each side=4x-1 square units

OpenStudy (anonymous):

1. b. a 2 −b 2 =(a+b)(a−b)

OpenStudy (anonymous):

c.) x 2 (x−9)−4(x−9)=(x−9)(x 2 −4) make also the factors of x 2 −4

OpenStudy (hoblos):

great job! :)

OpenStudy (anonymous):

thanks :)